# StereoWorld: Geometry-Aware Monocular-to-Stereo Video Generation

### v1

@@ Abstract
The growing adoption of XR devices has fueled strong demand for high-quality stereo video, yet its production remains costly and artifact-prone. To address this challenge, we present StereoWorld, an end-to-end framework that repurposes a pretrained video generator for high-fidelity monocular-to-stereo video generation. Our framework jointly conditions the model on the monocular video input while explicitly supervising the generation with a geometry-aware regularization to ensure 3D structural fidelity. A spatio-temporal tiling scheme is further integrated to enable efficient, high-resolution synthesis. To enable large-scale training and evaluation, we curate a high-definition stereo video dataset containing over 11M frames aligned to natural human interpupillary distance (IPD). Extensive experiments demonstrate that StereoWorld substantially outperforms prior methods, generating stereo videos with superior visual fidelity and geometric consistency. The project webpage is available at StereoWorld-Page.
@@ 1 Introduction
An alternative and more recent paradigm is the depth-warping-inpainting pipeline, powered by advances in diffusion models [dai2024svg, shi2024stereocrafter, wang2024stereodiffusion, zhao2024stereocrafter, zhang2024spatialme]. Depth is first estimated from the input video and used to warp frames into the target viewpoint; occluded regions are then hallucinated by an inpainting model. While conceptually simple, this pipeline suffers from distinct drawbacks. Specifically, the inpainting phase is decoupled from stereo geometry estimation, breaking pixel-level correspondence and resulting in texture distortions, color shifts, and stereo artifacts that degrade viewing comfort.
@@
  • Extensive experiments demonstrate that StereoWorld substantially outperforms prior works in visual quality, geometric consistency, and temporal stability, with clear advantages in both objective metrics and subjective perception.
@@ 2.1 Novel View Synthesis
With the advent of deep learning, novel view synthesis (NVS) has progressed rapidly, evolving from traditional Structure-from-Motion (SfM) [schonberger2016structure] methods to neural approaches such as NeRF [mildenhall2021nerf] and 3D Gaussian Splatting (3DGS) [kerbl20233d]. While NeRF implicitly models scene geometry, 3DGS introduces explicit representations that improve reconstruction quality and efficiency. Recent works extend NVS to dynamic scenes and feed-forward reconstruction. For example, 4DGS [wu20244d] and Shape of Motion [wang2024shape] reconstruct dynamic geometry from monocular videos, while methods such as MVSplat [chen2024mvsplat], PixelSplat [charatan2024pixelsplat], and VGGT [wang2025vggt] leverage large pretrained models for fast, pose-free reconstruction. However, reconstructed geometry remains sparse, and the visual fidelity of novel views is still limited, restricting their applicability to stereo video generation.
@@ 2.2 Diffusion-based Stereo Generation
Existing diffusion-based stereo generation methods—whether training-free approaches such as SVG [dai2024svg], StereoCrafter-Zero [shi2024stereocrafter], T-SVG [jin2024t], and StereoDiffusion [wang2024stereodiffusion], or pretrained models including StereoCrafter [zhao2024stereocrafter], SpatialMe [zhang2024spatialme], StereoConversion [mehl2024stereo], SpatialDreamer [lv2025spatialdreamer], ImmersePro [shi2024immersepro], ReStereo [huang2025restereo], and GenStereo [qiao2025genstereo]—generally follow a similar pipeline: monocular depth estimation, view warping, and diffusion-based inpainting of occluded regions. However, this paradigm disrupts the natural video distribution, causing spatial–temporal inconsistencies and degraded fidelity. In contrast, our method departs fundamentally from this paradigm by directly generating stereo videos in an end-to-end manner, thereby ensuring cross-view consistency and preserving visual fidelity.
@@ 3.3 Framework
The essence of stereoscopy lies in the depth variations among scene objects, which define their hierarchical spatial relationships relative to the observer. However, relying solely on monocular conditioning and a standard $\mathcal{L}_{\text{rgb}}$ reconstruction loss is insufficient to capture such geometric structure. As shown in Fig. 8, we observe that the model struggles to implicitly learn this complex geometric structure from only RGB data, often resulting in outputs with weak stereo perception, such as flattened object boundaries or unstable disparities. This suggests that the model requires a more explicit signal to guide its understanding of 3D geometry. Therefore, we propose a geometry-aware regularization strategy to enhance the model’s 3D perception capacity.
@@
Temporal Tiling Strategy. Our base model generates only short clips (81 frames, about 3s at 24 FPS). To handle longer videos, we split them into overlapping segments, using the last frames of each segment to guide the next for smooth transitions [zhao2024stereocrafter]. To further reduce flickering, during training we probabilistically replace the first few frames of noisy latents with clean frames (probability $p$), enabling the model to learn robust long-range temporal consistency (Fig. 4).

### v2

@@ Abstract
The growing adoption of XR devices has fueled strong demand for high-quality stereo video, yet its production remains costly and artifact-prone. To address this challenge, we present StereoWorld, an end-to-end framework that repurposes a pretrained video generator for high-fidelity monocular-to-stereo video generation. Our framework jointly conditions the model on the monocular video input while explicitly supervising the generation with a geometry-aware regularization to ensure 3D structural fidelity. A spatio-temporal tiling scheme is further integrated to enable efficient, high-resolution synthesis. To enable large-scale training and evaluation, we curate a high-definition stereo video dataset containing over 11M frames aligned to natural human interpupillary distance (IPD). Extensive experiments demonstrate that StereoWorld substantially outperforms prior methods, generating stereo videos with superior visual fidelity and geometric consistency.
@@ 1 Introduction
An alternative and more recent paradigm is the depth-warping-inpainting pipeline, powered by advances in diffusion models [dai2024svg, shi2024stereocrafter, wang2024stereodiffusion, zhao2024stereocrafter, zhang2024spatialme]. Depth is first estimated from the input video and used to warp frames into the target viewpoint; occluded regions are then hallucinated by an inpainting model to plausibly fill in the missing areas. While conceptually simple, this pipeline suffers from distinct drawbacks. Specifically, the inpainting phase is decoupled from stereo geometry estimation, breaking pixel-level correspondence and resulting in texture distortions, color shifts, and stereo artifacts that degrade viewing comfort.
@@
  • Extensive experiments demonstrate that StereoWorld substantially outperforms prior works in visual quality, geometric consistency, and temporal stability, with clear advantages in objective metrics and subjective perception.
@@ 2.1 Novel View Synthesis
With the advent of deep learning, novel view synthesis (NVS) has progressed rapidly, evolving from traditional Structure-from-Motion (SfM) [schonberger2016structure] methods to neural approaches such as NeRF [mildenhall2021nerf] and 3D Gaussian Splatting (3DGS) [kerbl20233d]. While NeRF implicitly models scene geometry, 3DGS introduces explicit representations that improve reconstruction quality and efficiency. Recent works extend NVS to dynamic scenes and feed-forward reconstruction. For example, 4DGS [wu20244d] and Shape of Motion [wang2024shape] reconstruct dynamic geometry from monocular videos, while methods like MVSplat [chen2024mvsplat], PixelSplat [charatan2024pixelsplat], and VGGT [wang2025vggt] leverage large pretrained models for fast, pose-free reconstruction. However, reconstructed geometry remains sparse, and the visual fidelity of novel views is still limited, restricting their applicability to stereo video generation.
@@ 2.2 Diffusion-based Stereo Generation
Existing diffusion-based stereo generation methods—whether training-free approaches such as SVG [dai2024svg], StereoCrafter-Zero [shi2024stereocrafter], T-SVG [jin2024t], and StereoDiffusion [wang2024stereodiffusion], or pretrained models including StereoCrafter [zhao2024stereocrafter], SpatialMe [zhang2024spatialme], StereoConversion [mehl2024stereo], SpatialDreamer [lv2025spatialdreamer], ImmersePro [shi2024immersepro], ReStereo [huang2025restereo], and GenStereo [qiao2025genstereo]—generally follow a similar pipeline: monocular depth estimation, view warping, and diffusion-based inpainting of occluded regions. But this paradigm disrupts the natural video distribution, causing spatial–temporal inconsistencies and degraded fidelity. In contrast, our method departs fundamentally from this paradigm by directly generating stereo videos in an end-to-end manner, thereby ensuring cross-view consistency and preserving visual fidelity.
@@ 3.3 Framework
The essence of stereoscopy lies in the depth variations among scene objects, which define their hierarchical spatial relationships relative to the observer. However, relying solely on monocular conditioning and a standard $\mathcal{L}_{\text{rgb}}$ reconstruction loss is insufficient to capture such geometric structure. As shown in Fig. 8, we observe that the model struggles to implicitly learn complex geometric structures from only RGB data, resulting in outputs with weak stereo perception, such as flattened object boundaries or unstable disparities. This suggests that the model requires a more explicit signal to guide its understanding of 3D geometry. Therefore, we propose a geometry-aware regularization strategy to enhance the model’s 3D perception capacity.
@@
Temporal Tiling Strategy. Our base model generates only short clips (81 frames, about 3s at 24 FPS). To handle longer videos, we split them into overlapping segments, using the last frames of each segment to guide the next for smooth transitions [zhao2024stereocrafter]. To further reduce flickering, during training we probabilistically replace first few frames of noisy latents with clean frames (probability $p$), enabling the model to learn robust long-range temporal consistency (Fig. 4).